Consider the following situation׃ the force exerted on the small piston of a hydraulic lift is 160 N. The areas of the small piston and the large piston are 0.5 m2 and 1.0 m2, respectively. 


1.     What is the pressure exerted on the large piston?

A.   80 Pa                                                                        C. 320 Pa

B.    160 Pa                                                                      D. 640 Pa

 

2.     What is the output force on the large piston?

A.   80 N                                                                          C. 320 N

B.    160 N                                                                        D. 640 N

 

3.     Which term in Bernoulli’s theorem is referred to as the pressure head?

A.   mgh    C. v2/2g

B.     P𝑚/𝜌     D.    P/pg

1
Expert's answer
2022-02-07T15:27:16-0500

1.


"F_2=F_1\frac{A_2}{A_1}=\rm160\: N*\frac{1.0\: m^2}{0.5\: m^2}=320\: N"

Answer: C


2.

Answer: C


3.

Answer: D